OpenStudy (dls):

Trout are bred in a tank as fish for restaurants Twenty trout are caught and each is tagged through the tail then released back in the tank A week later,30 trout are caught and it is found that two of them have a tail tag Based on this evidence alone,what estimate can be made about no. of trout in the tank 100 300 600 1200

Directrix (directrix):

20 trout are tagged from of the tank containing an unknown number of trout. 30 trout are captured from the tank. Two of these 30 are tagged. 2 is to 30 as 20 is to T where T is the estimated number of fish in the tank.

Directrix (directrix):

2/30 = 20/T 1/15 = 20/ T Your task: Solve the above equation for T.

Directrix (directrix):

T = ?

OpenStudy (dls):

300

Directrix (directrix):

That's what I got.

OpenStudy (dls):

thanks!

Directrix (directrix):

Glad to help.